find the surface area of the object obtained by rotating y=4+3x^2, 1<=x<=2 about the y axis


1
Expert's answer
2022-04-08T13:32:25-0400

"y=4+3x^2" , "1\leq x\leq2"

Express x as function of y:

"x=\frac{1}{\sqrt3}\sqrt{y-4}" , "7\leq y\leq 16"



The surface area can be calculated using the formula:

"S=2\pi\int_7^{16} x(y)\sqrt{1+(x'(y))^2}dy"

"x'=\frac{1}{2\sqrt3}\frac{1}{\sqrt{y-4}}"

"S=2\pi\int_7^{16} \frac{1}{\sqrt3}\sqrt{y-4}\sqrt{1+\frac{1}{12}\frac{1}{y-4}}dy=""2\pi\int_7^{16} \frac{1}{\sqrt3}\sqrt{y-4}\sqrt{1+\frac{1}{12}\frac{1}{y-4}}dy=""\frac{1}{3}\pi\int_7^{16} \sqrt{12y-47}dy=\frac{1}{3}\pi\cdot\frac{2}{3\cdot12}(12y-47)^{3/2}|_7^{16}=""\frac{1}{3}\pi\cdot\frac{2}{3\cdot12}(12y-47)^{3/2}|_7^{16}=\frac{\pi}{54}(145^{3/2}-37^{3/2})""\approx88.49"